OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. cos 94° cos 37° + sin 94° sin 37°

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@RadEn

OpenStudy (dumbcow):

\[\cos(a-b) = \cos a \cos b + \sin a \sin b\]

OpenStudy (anonymous):

sin57

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

um no, did you misread the identity i posted check it again

OpenStudy (anonymous):

ohh cos57

OpenStudy (dumbcow):

haha yep :)